OpenStudy (anonymous):

(its a sigma notation )Given the arithmetic sequence 23,17,11...evaluate

OpenStudy (anonymous):

\[\sum_{n=4}^{9} a _{n}\]

OpenStudy (amistre64):

what our common difference between terms?

OpenStudy (amistre64):

23 + c = 17 17+c = 11 what is c?

OpenStudy (amistre64):

once you know c you can determine the values of a4 and a9 then add them up 5 times and divide in half

OpenStudy (anonymous):

i'm still confused

OpenStudy (anonymous):

where are you getting c from?

OpenStudy (amistre64):

an arithmetic sequence is created by adding a constant value to the last term to get the next term. all you need to do to find the common term is to work out the addition from one term to the next

OpenStudy (amistre64):

our 1st terms is 23 ; they add some amount to it to get 17 23 + c = 17

OpenStudy (anonymous):

so c is constant?

OpenStudy (amistre64):

yes, also known as the common difference between each term since its value never changes

OpenStudy (anonymous):

so 6 ?

OpenStudy (amistre64):

well, -6 in this case :)

OpenStudy (anonymous):

23+-6=17 17+-6=11

OpenStudy (amistre64):

now with that iinformation we can either construct the sequence and add up the 4th to 9th terms

OpenStudy (anonymous):

but what do i do next i dnt get the\[\sum_{?}^{?}\] thing?

OpenStudy (amistre64):

or do some fancier stuff that does the same thing

OpenStudy (anonymous):

ignore the questions marks please

OpenStudy (anonymous):

so the answer is -6?

OpenStudy (amistre64):

the 4 on the bottom means we start with the 4th term and add up all the terms till we get to the 9th term; since we got a 9 up top

OpenStudy (anonymous):

how do i evaluate E

OpenStudy (amistre64):

what are the 4th 5th 6th 7th 8th and 9th terms?

OpenStudy (anonymous):

what do you mean 4th terms?

OpenStudy (amistre64):

E is just a notation that tells us we are summing up (adding) all the terms from the bottom-th to the top-th

OpenStudy (amistre64):

you see that 23 in front of our sequence?

OpenStudy (anonymous):

so 9+4?

OpenStudy (anonymous):

yeah...

OpenStudy (amistre64):

the 23 is our 1st term in the sequence do you see that 17 that is right next to it?

OpenStudy (anonymous):

yeah so that;s the 2nd and 11 is the 3rd?

OpenStudy (amistre64):

yep; and we know how to find each next term so we just need to find the 4th to the 9th terms

OpenStudy (anonymous):

so 4th terms is 5?

OpenStudy (anonymous):

and 5th is -1

OpenStudy (anonymous):

and so on?

OpenStudy (amistre64):

 yep

OpenStudy (anonymous):

so what do i do next after i have all terms 4th-9th?

OpenStudy (amistre64):

term: 4 5 6 7 8 9 value: 5 -1 -7 -13 -19 -25 add em all up

OpenStudy (anonymous):

so -60?